# A Multi-Task Foundation Model for Wireless Channel Representation Using Contrastive and Masked Autoencoder Learning

### v1

@@ IV-D1 Cross-frequency Beam Selection
Figure 4: Top-1 beam selection accuracy for a codebook size of 32 achieved by WiMAE/ContraWiMAE vs. baselines followed by (a) linear probing, (b) ResNet-Slim, and (c) ResNet-Wide.
TABLE III: Top-1/top-3 accuracy for beam selection with linear probing.
Model Codebook Size (CS) Training Budget
1% 2% 5% 10% 25% 50% 100%
WiMAE 16 54.8/77.1 $62.7$/$82.8$ $67.0$/$85.2$ $73.1$/$86.1$ $79.7$/$90.6$ $81.4$/$91.4$ $79.2$/$87.1$
32 39.9/58.2 $47.2$/$66.0$ $55.3$/$74.6$ $61.7$/$74.9$ $72.9$/$82.5$ $74.8$/$81.0$ $72.4$/$76.6$
64 $23.6$/$40.8$ $31.1$/$50.9$ $30.7$/$45.4$ $40.8$/$57.0$ $55.3$/$68.9$ $57.8$/$66.8$ $64.1$/$72.9$
128 $12.7$/$22.0$ $16.9$/$28.0$ $17.5$/$30.2$ $23.6$/$35.4$ $31.5$/$45.3$ $35.6$/$47.6$ $39.8$/$46.6$
256 $6.8$/$12.5$ $10.4$/19.0 $11.7$/$19.4$ $12.7$/$20.0$ $17.6$/$26.6$ $22.9$/$31.5$ $23.3$/$31.7$
ContraWiMAE 16 $54.2$/$76.4$ 65.3/85.9 73.8/89.1 76.2/87.9 81.3/92.5 83.0/91.6 85.0/92.6
32 $39.0$/$55.2$ 53.0/74.5 64.8/81.2 77.0/88.6 84.0/94.3 85.4/92.3 85.6/92.7
64 24.9/41.2 32.5/51.8 44.7/66.7 55.0/74.6 67.7/84.4 75.8/88.3 74.7/82.0
128 15.0/24.9 18.3/32.4 25.0/41.0 28.8/46.6 42.5/62.8 51.4/70.2 55.9/70.6
256 7.9/13.5 10.5/$18.7$ 14.7/24.1 17.3/27.2 22.5/35.4 27.4/41.5 31.3/45.2
LWM 16 $17.5$/$27.5$ $23.6$/$35.3$ $19.0$/$32.2$ $26.0$/$36.9$ $36.8$/$48.5$ $39.6$/$47.9$ $42.7$/$53.6$
32 $18.8$/$27.3$ $19.0$/$26.1$ $16.3$/$22.5$ $26.6$/$34.3$ $34.2$/$41.0$ $40.1$/$47.2$ $43.3$/$49.1$
64 $14.8$/$20.7$ $18.9$/$25.8$ $11.8$/$16.4$ $21.3$/$27.2$ $27.8$/$35.0$ $33.0$/$39.7$ $36.7$/$42.6$
128 $10.0$/$13.5$ $12.0$/$19.1$ $8.2$/$13.1$ $14.1$/$20.5$ $19.2$/$25.5$ $24.1$/$29.7$ $25.1$/$29.2$
256 $5.3$/$8.0$ $8.6$/$12.9$ $6.0$/$10.4$ $8.3$/$12.2$ $11.6$/$17.1$ $13.7$/$18.8$ $13.7$/$19.2$
RAW 16 $31.6$/$41.5$ $33.9$/$42.8$ $34.0$/$43.7$ $38.2$/$45.8$ $38.8$/$46.3$ $39.8$/$44.4$ $41.7$/$48.0$
32 $28.5$/$38.9$ $27.6$/$37.5$ $37.1$/$45.8$ $37.6$/$45.0$ $37.5$/$44.7$ $41.2$/$47.5$ $42.6$/$47.6$
64 $19.4$/$29.2$ $25.2$/$36.2$ $29.3$/$41.6$ $32.0$/$42.8$ $34.6$/$43.1$ $37.1$/$44.5$ $36.5$/$42.2$
128 $10.9$/$19.3$ $15.0$/$25.3$ $20.7$/$32.1$ $24.0$/$35.0$ $25.3$/$36.8$ $29.6$/$40.6$ $30.7$/$38.0$
256 $6.2$/$10.9$ $9.9$/$15.8$ $11.5$/$19.1$ $15.0$/$24.9$ $18.7$/$26.1$ $20.0$/$31.5$ $22.7$/$31.7$
Note: Boldface indicates the best result for each (CS, training budget) pair.
TABLE IV: Top-1/top-3 accuracy for beam selection with ResNet-Slim and ResNet-Wide.
Model Codebook Size (CS) Training Budget
ResNet-Slim ResNet-Wide
10% 25% 50% 100% 10% 25% 50% 100%
WiMAE 16 33.0/60.9 67.8/86.2 74.0/90.2 79.9/93.0 $\underline{\textbf{64.3}}$/$\underline{\textbf{83.2}}$ $\underline{\textbf{78.6}}$/$\underline{\textbf{92.6}}$ $\underline{\textbf{81.9}}$/$\underline{\textbf{94.7}}$ $\underline{\textbf{84.0}}$/$\underline{\textbf{96.0}}$
32 22.6/46.6 65.1/88.9 76.0/$93.3$ $81.9$/$95.5$ $\underline{\textbf{54.5}}$/$\underline{\textbf{84.7}}$ $78.4$/$95.0$ $\underline{\textbf{85.9}}$/$97.2$ $87.5$/$97.4$
64 $2.8$/$7.9$ 32.8/61.7 63.9/90.6 $76.0$/$95.3$ $\underline{\textbf{28.1}}$/$\underline{\textbf{58.0}}$ $\underline{\textbf{64.7}}$/$90.6$ $74.6$/$95.3$ $\underline{\textbf{81.0}}$/$97.2$
128 6.1/13.1 15.8/35.4 36.5/68.7 $55.8$/$88.1$ $\underline{\textbf{14.0}}$/$\underline{\textbf{30.0}}$ $\underline{\textbf{38.4}}$/$\underline{\textbf{71.4}}$ $\underline{\textbf{59.4}}$/$89.5$ $\underline{\textbf{70.0}}$/$93.5$
256 $1.1$/$2.9$ 8.0/16.4 $17.3$/40.0 37.7/73.0 $\underline{\textbf{8.0}}$/$\underline{\textbf{15.8}}$ $\underline{\textbf{18.2}}$/$\underline{\textbf{40.2}}$ $30.8$/$66.0$ $\underline{\textbf{49.9}}$/$85.6$
ContraWiMAE 16 $30.9$/$55.5$ $59.8$/$80.5$ $73.3$/$90.2$ 79.9/$92.6$ $58.2$/$81.7$ $78.0$/$\underline{\textbf{92.6}}$ $80.7$/$94.4$ $83.5$/$95.2$
32 $17.9$/$39.8$ $46.4$/$74.8$ $74.7$/94.1 84.5/96.4 $47.3$/$71.6$ $\underline{\textbf{79.7}}$/$\underline{\textbf{96.0}}$ $85.7$/$\underline{\textbf{97.3}}$ $\underline{\textbf{87.7}}$/$\underline{\textbf{97.5}}$
64 11.3/26.8 $28.0$/$58.4$ $60.4$/$89.4$ 76.8/96.4 $24.5$/$50.0$ $62.5$/$\underline{\textbf{90.7}}$ $\underline{\textbf{75.6}}$/$\underline{\textbf{96.8}}$ $80.9$/$\underline{\textbf{98.0}}$
128 $5.0$/$12.1$ $14.5$/$31.1$ $34.0$/$65.0$ 60.2/91.9 $13.6$/$27.6$ $33.3$/$64.6$ $59.2$/$\underline{\textbf{92.3}}$ $69.3$/$\underline{\textbf{95.6}}$
256 3.6/6.4 $7.7$/$15.6$ 18.3/$38.8$ $36.9$/$72.6$ $6.7$/$13.8$ $14.9$/$32.0$ $\underline{\textbf{34.4}}$/$\underline{\textbf{71.7}}$ $49.7$/$\underline{\textbf{87.5}}$
LWM 16 $23.0$/$54.5$ $47.0$/$81.7$ $61.5$/$87.6$ $69.2$/$91.0$ $49.5$/$80.3$ $70.0$/$91.0$ $74.4$/$92.5$ $79.6$/$95.7$
32 $12.3$/$32.6$ $33.1$/$67.1$ $55.3$/$86.1$ $71.3$/$94.2$ $33.8$/$65.0$ $67.3$/$90.1$ $76.8$/$97.0$ $81.3$/$96.8$
64 $2.5$/$5.6$ $15.8$/$40.9$ $36.0$/$71.3$ $53.5$/$89.0$ $18.9$/$40.6$ $36.5$/$70.2$ $59.4$/$91.3$ $71.4$/$97.1$
128 $0.7$/$1.9$ $6.8$/$16.1$ $13.7$/$35.3$ $36.9$/$74.9$ $7.8$/$18.6$ $19.6$/$41.8$ $34.6$/$71.0$ $54.6$/$90.6$
256 $0.2$/$1.1$ $5.1$/$9.9$ $7.8$/$16.2$ $19.6$/$48.4$ $4.7$/$10.6$ $7.8$/$18.4$ $14.6$/$35.4$ $29.0$/$63.9$
RAW 16 $7.9$/$25.7$ $31.4$/$58.5$ $64.1$/$81.8$ $70.6$/$88.5$ $28.7$/$52.1$ $54.3$/$78.4$ $73.0$/$88.9$ $79.3$/$92.7$
32 $4.0$/$13.1$ $27.2$/$54.7$ $55.4$/$79.0$ $72.7$/$91.6$ $24.3$/$43.6$ $50.0$/$76.3$ $72.9$/$92.9$ $86.1$/$97.0$
64 2.8/$5.1$ $16.9$/$32.6$ $29.6$/$53.2$ $55.0$/$80.7$ $15.7$/$30.3$ $37.3$/$63.1$ $59.3$/$82.8$ $77.6$/$94.9$
128 $1.6$/$4.3$ $8.1$/$18.4$ $20.6$/$38.8$ $45.0$/$76.4$ $8.7$/$17.9$ $24.0$/$43.6$ $44.8$/$72.8$ $66.4$/$90.8$
256 $0.7$/$1.4$ $0.2$/$1.9$ $10.2$/$22.0$ $24.7$/$49.5$ $0.9$/$1.4$ $12.4$/$24.3$ $23.4$/$46.8$ $47.1$/$78.5$
IV-D2 LoS Detection
This task classifies if a channel exhibits the LoS condition, which can be crucial for adaptive modulation and coding, blockage prediction, and predictive handover. We report accuracy, F1 score, and area under the ROC curve (AUC), comparing the WiMAE and ContraWiMAE representations with the LWM baseline. In our comparison, we only employ linear probing since complex downstream models offer minimal additional improvement for this task. We use the CLS token for LWM and mean-pooled embeddings for WiMAE and ContraWiMAE. Although WiMAE and ContraWiMAE do not have a classification token like CLS, calculating the mean value across each patch encoding serves as a summary of the input. All experiments, shown in Table V, use Adam optimizer with a learning rate of 0.01 with exponential decay with $\gamma=0.995$ and a batch size of 256.
TABLE V: Accuracy in the LoS detection task.
Training Budget (%) LWM WiMAE ContraWiMAE
Accuracy F1 AUC Accuracy F1 AUC Accuracy F1 AUC
1 $0.929$ $0.941$ 0.981 $0.931$ 0.944 $0.980$ 0.931 0.944 0.981
2 0.946 0.957 0.985 $0.941$ $0.953$ $0.983$ $0.941$ $0.953$ $0.983$
5 $0.926$ $0.942$ $0.973$ $0.944$ $0.955$ $0.985$ 0.949 0.959 0.987
10 $0.943$ $0.955$ $0.987$ $0.947$ $0.958$ 0.989 0.950 0.960 0.989
25 $0.946$ $0.957$ $0.988$ 0.957 0.965 0.992 $0.951$ $0.960$ $0.990$
50 $0.950$ $0.960$ $0.989$ 0.958 0.966 0.992 $0.956$ $0.965$ 0.992
100 $0.948$ $0.959$ $0.989$ 0.960 0.968 0.993 $0.959$ $0.967$ $0.992$
Note: Boldface indicates best result for each training budget.
V Analysis of Results
In this section, we examine in detail various model configurations and performance metrics across different downstream tasks and extract insights into optimal architectural design, pretraining strategies, and fundamental relationships between model capacity, task complexity, and data efficiency. Our comprehensive evaluation reveals both the strengths of our approaches and the nuanced interplay between different architectural choices.
V-A Optimal Architectural Design of WiMAE
We begin by discussing the performance of WiMAE on the beam prediction downstream task, analyzing the impact of the masking ratio, the model depth, and the representation dimension. We also provide insights on the masked channel reconstruction, which is at the base of WiMAE’s pretraining stage.
@@ V-A Optimal Architectural Design of WiMAE
Fig. 2(a) reveals that for all codebook sizes considered, WiMAE’s performance peaks at a masking ratio of $m_{\text{r}}=0.6$, with a notable decline at higher ratios. This suggests an optimal information density for these tasks where sufficient masking forces the model to learn meaningful representations without compromising its ability to capture essential channel characteristics. At lower masking ratios (0.45), the pretraining task becomes less challenging, potentially leading to over-reliance on local correlations rather than learning global structural properties. In contrast, at extreme masking ratios (0.9), the reconstruction task approaches infeasibility, forcing the model to resort to generic patterns rather than capturing task-specific features.
@@
Fig. 2(b) demonstrates the complex relationship between model depth and performance. As the encoder depth increases from 6 to 12 layers, we observe consistent performance improvements across all codebook sizes. However, the marginal benefit diminishes beyond 12 layers, with 18-layer models showing minimal improvement. This suggests that moderate encoder depth effectively captures channel characteristics without excessive computation or potential overfitting. More interestingly, for a fixed 12-layer encoder, increasing the decoder depth from 4 to 8 and then to 12 layers provides substantial performance improvements. Since the decoder is discarded after pretraining, its added complexity does not affect inference latency. This asymmetric architecture—moderate encoder with deep decoder—reveals that a sophisticated decoding process provides a stronger learning signal that helps the encoder capture more nuanced channel features during training.
@@
Fig. 2(c) reveals a strong positive correlation between embedding dimension and performance, with patterns that vary by task complexity. Top-1 accuracy improves as the dimension increases from 32 to 128, with diminishing returns becoming apparent beyond 96. The most significant improvement occurs between dimensions 32 and 64, suggesting that this range is critical for adequately representing channel characteristics for the data under consideration. More challenging tasks show less saturation at higher dimensions, indicating that complex discrimination tasks benefit more from increased representational capacity.
Reconstruction performance
Fig. 3 shows the reconstruction of both magnitude and phase information from heavily masked inputs using settings described in Section IV-B. As shown in the figure, WiMAE demonstrates impressive capability in capturing structural patterns rather than simply interpolating between known values. The reconstructions preserve characteristic fading patterns, spatial correlations across antenna elements, and coherent phase relationships between antenna elements and subcarriers. The transitions between visible and reconstructed regions are smooth, without boundary artifacts, demonstrating spatial coherence and the model’s ability to leverage global context. This quality of reconstruction with 60% masking suggests that WiMAE learns meaningful representations of the underlying channel physics rather than memorizing patterns, making it well-suited for fine-tuning on task-specific scenarios.
V-B Comparison of WiMAE, ContraWiMAE, and Baselines
In this section, we compare the performance of WiMAE and ContraWiMAE to other baselines for the beam prediction and LoS detection downstream tasks. The full evaluation is reported in Table III, Table IV, and Table V, whereas Fig. 4 highlights the accuracy vs. training budget tradeoff obtained on beam selection for a codebook size of 32 and different downstream models.
Linear probing on beam selection
The results in Fig. 4(a) demonstrate WiMAE’s data efficiency. For codebook size 32, with only 1% training data, WiMAE achieves 39.9% top-1 accuracy, surpassing both LWM (18.8%) and raw models (28.5%). Notably, WiMAE with 5% training data outperforms LWM trained on the entire dataset, demonstrating superior feature extraction capabilities. The addition of contrastive learning in ContraWiMAE substantially enhances the linear separability of the encoded representations. With linear probing, as the training budget increases from 1% to 100%, ContraWiMAE demonstrates consistent performance scaling, with the advantage over WiMAE growing progressively larger. From Table III, we observe that this scaling effect becomes more pronounced with complex tasks (i.e., larger beam codebook sizes), where the contrastive objective promotes a more homogeneous utilization of the latent space, enabling clearer decision boundaries even as class complexity increases. This pattern suggests that contrastive learning enables the model to maintain discriminative power even as classification complexity increases, creating more linearly separable embeddings that preserve class distinctions without requiring complex downstream architectures.
ResNet models on beam selection
Figs. 4(b) and 4(c) examine the top-1 beam selection accuracy with a codebook size of 32 when the downstream model is a ResNet-Slim and a ResNet-Wide, respectively. While ContraWiMAE outperforms LWM by 42.3% with linear probing, this gap narrows to just 6.4% with ResNet-Wide, demonstrating how complex downstream architectures can offset weaker foundation representations. Table IV further shows that WiMAE and ContraWiMAE maintain their edge across configurations, especially in data-constrained scenarios. Notably, for a codebook size of 16 and 100% of the training budget, upgrading from linear probing to ResNet-Wide improves the top-1 accuracy for WiMAE by only 4.8%, while it improves LWM by 36.9%. This indicates that WiMAE’s representations are already well optimized. For complex tasks, increased downstream capacity benefits all models, though WiMAE maintains higher absolute performance.
Linear probing on LoS detection
Table V examines the performance of different models for LoS detection. Both WiMAE and LWM exceed 92% accuracy even with just 1% training data, suggesting that LoS characteristics are more easily extractable compared to beam indices. WiMAE maintains a slight advantage with 100% training data (96.0% accuracy vs. 94.8% for LWM). WiMAE shows consistent improvement as the training data increases, while LWM exhibits small fluctuations, suggesting that WiMAE provides more stable feature spaces. ContraWiMAE shows competitive performance with WiMAE, while consistently outperforming LWM, suggesting that introducing the contrastive objective does not hurt the encoder’s ability to produce meaningful compact summary representations.
VI Conclusion
In this paper, we introduced WiMAE, a transformer-based masked autoencoder specifically designed for wireless channel representation learning. By employing an asymmetric encoder-decoder architecture, high masking ratios, and moderate encoder depth, WiMAE strikes a balance between representational quality and computational efficiency. Our experiments showed that WiMAE consistently outperforms existing baselines across multiple downstream tasks, achieving particularly strong results under constrained data conditions. This efficiency is especially critical for practical wireless scenarios where labeled data is expensive or difficult to obtain.
To further enhance the quality of learned representations, we proposed ContraWiMAE, which extended WiMAE by incorporating a contrastive learning objective in a unified multi-task framework. Warm-starting from pretrained WiMAE weights and using additive noise to generate positive pairs, ContraWiMAE captures both structural and discriminative features of wireless channels. This hybrid approach further improves the linear separability of the learned embeddings, leading to enhanced downstream performance, especially with simple classifiers. Importantly, ContraWiMAE preserves the strong reconstruction performance of its base model, demonstrating that the two learning objectives can effectively complement each other.
Through extensive evaluation on realistic, previously unseen wireless scenarios, both WiMAE and ContraWiMAE exhibited strong transferability and robustness. WiMAE achieved optimal performance at a masking ratio of 0.6, with deeper decoders offering better representations without affecting inference latency. Larger embedding dimensions were found to be especially beneficial for complex tasks, and the data efficiency of both models enables high performance using as little as 1% of the training data employed by other baselines. Unlike other models that require complex downstream architectures to achieve good performance, WiMAE and ContraWiMAE produce representations that remain effective even with lightweight classifiers.

### v2

@@ VI-A Comparison of WiMAE, ContraWiMAE, and Baselines
Figure 3: Top-1 beam selection accuracy for a codebook size of 32 achieved by ContraWiMAE vs. baselines followed by (a) linear probing, (b) ResNet-Slim, and (c) ResNet-Wide.
In this section, we compare the performance of ContraWiMAE with those of WiMAE and other baselines for beam selection (Tables II, III), LoS detection (Table IV), and channel estimation (Fig. 2) downstream tasks. In addition, Fig. 3 highlights the accuracy vs. training budget tradeoff obtained on beam selection for a codebook size of 32 and different downstream models. For the results in this section, we use a vertical patch shape with $(N_{\text{p,s}},N_{\text{p,f}})=(16,1)$ and $m_{\text{r}}=0.9$ as shown in Fig. 4, since it accommodates the channel estimation task where we consider a pilot tone at a single subcarrier to estimate the channel.
Figure 4: Magnitude (left) and phase (right) reconstruction for ContraWiMAE, obtained with patch size $(N_{\text{p,s}},N_{\text{p,f}})={\color[rgb]{0,0,0}(16,1)}$. Each row is a channel sample. Columns correspond to original, masked, and reconstructed channels.
Linear probing on beam selection
The results in Fig. 3(a) demonstrate ContraWiMAE’s data efficiency. For codebook size 32, with only 1% training data, ContraWiMAE achieves 35.9% top-1 accuracy, surpassing all other baselines. Notably, ContraWiMAE with 2% training data outperforms LWM trained on the entire dataset, demonstrating superior feature extraction capabilities. The addition of masked contrastive component in ContraWiMAE over WiMAE substantially enhances the linear separability of the encoded representations, highlighting that combining contrastive learning with masked autoencoding addresses fundamental limitations of WiMAE for wireless data. With linear probing, as the training budget increases from 1% to 100%, ContraWiMAE demonstrates consistent performance scaling, with the advantage over other baselines growing progressively larger. This pattern suggests that contrastive learning enables the model to gain substantial discriminative power, creating more linearly separable embeddings that preserve class distinctions without requiring complex downstream architectures.
ResNet models on beam selection
Figs. 3(b) and 3(c) examine the top-1 beam selection accuracy with a codebook size of 32 when the downstream model is a ResNet-Slim and a ResNet-Wide, respectively. While ContraWiMAE outperforms LWM by 35.7% with linear probing, this gap narrows to just 4.9% with ResNet-Wide, demonstrating how complex downstream architectures can offset weaker foundation representations. The diminishing gains with complex models like ResNet-Wide demonstrate the effectiveness of our pretraining. Complex downstream models can compensate for poor initial representations when given sufficient task-specific data, while our pretrained representations are already well-suited for the tasks, resulting in consistently strong performance without requiring extensive downstream model complexity. Table III further shows that ContraWiMAE maintains its edge across all configurations. Notably, for small codebooks at 100% training budget, upgrading from linear probing to ResNet-Wide leads to minimal performance gain for ContraWiMAE, while for larger codebooks the difference becomes more pronounced. This indicates that ContraWiMAE’s representations are already well optimized for simpler tasks. For complex tasks, increased downstream capacity benefits all models.
Linear probing on LoS detection
Table IV examines the performance of different models for LoS detection. All models exceed 92% accuracy even with just 1% training data, suggesting that LoS characteristics are more easily extractable compared to beam indices. WiMAE maintains a slight advantage over other models. ContraWiMAE shows competitive performance with WiMAE, while consistently outperforming LWM, suggesting that introducing the contrastive objective does not hurt the encoder’s ability to produce meaningful compact summary representations.
(a) In-distribution, Beam size 16
(b) Same-frequency, Beam size 16
(c) Cross-frequency, Beam size 16
(d) In-distribution, Beam size 32
(e) Same-frequency, Beam size 32
(f) Cross-frequency, Beam size 32
Figure 5: Beam selection performance under varying SNR for in-distribution, same-frequency OOD, and cross-frequency OOD settings with beam sizes 16 (top row) and 32 (bottom row)
Channel Estimation
Although ContraWiMAE and WiMAE are pretrained under different scenarios with randomly applied masks, as shown in Fig. 2(a), their performance remains comparable to the supervised model trained on the pretraining data with a single reconstruction mask that coincides with pilot positions. The dual objective of ContraWiMAE introduces only a slight reduction in reconstruction accuracy and its performance closely follows that of WiMAE. As expected, the supervised model trained with the same noise levels as those used during testing achieves the lowest NMSE, benefiting from its explicit exposure to noise characteristics during training.
Fig. 2(b) further illustrates ContraWiMAE’s scaling behavior with respect to the amount of finetuning data. The performance gains become less pronounced as the finetuning budget increases. Notably, when ContraWiMAE is finetuned on only 2.5% of the task dataset corresponding to around 7000 channels, it surpasses all the supervised baselines across all SNR levels, highlighting its superior sample efficiency.
Reconstruction performance
Fig. 4 shows the reconstruction of both magnitude and phase information from heavily masked inputs using settings described in Section IV-B. As shown in the figure, ContraWiMAE demonstrates impressive capability in capturing structural patterns rather than simply interpolating between known values. The reconstructions preserve characteristic fading patterns, spatial correlations across antenna elements, and coherent phase relationships between antenna elements and subcarriers. The transitions between visible and reconstructed regions are smooth, without boundary artifacts, demonstrating spatial coherence and the model’s ability to leverage global context. This quality of reconstruction with 90% masking suggests that ContraWiMAE learns meaningful representations of the underlying channel physics rather than memorizing patterns, making it well-suited for fine-tuning on task-specific scenarios.
VI-B Robustness Analysis
We study the ContraWiMAE’s robustness by examining the beam selection performance of a simple kNN classifier trained on encoder representations obtained under various distribution shift conditions as shown in Fig. 5. Under in-distribution evaluation, where the train and test points of kNN are from the same downstream task scenarios and the beam labels are also obtained at 3.5 GHz, we evaluate the noise robustness of ContraWiMAE. Figs. 5(a) and 5(d) show that ContraWiMAE progressively narrows the gap with supervised approaches. Remarkably, in the most challenging 0-10 dB SNR regime, our method achieves parity with supervised training. We next assess the generalization to out-of-distribution (OOD) scenarios through cross-scenario evaluation (Figs. 5(b) and 5(e)), where test environments remain completely separate from those used in kNN training. Here, ContraWiMAE retains its advantage over supervised baseline at lower SNR levels, with all methods exhibiting comparable degradation.
The most demanding evaluation appears in Figs 5(c) and 5(f), where we simultaneously introduce both unseen test scenarios and unseen carrier frequency of 28 GHz, creating both an environmental and spectral shift. Despite operating without labeled supervision, ContraWiMAE’s inherent noise robustness enables it to surpass the supervised baseline when these compounded shifts coincide with adverse noise conditions.
Figure 6: t-SNE representation with LoS (blue) and non-LoS (orange) labels under varying SNR
Additionally, we present t-SNE visualizations in Fig. 6 to illustrate the robustness of ContraWiMAE compared to WiMAE under varying signal-to-noise ratio conditions (0, 5, and 10 dB). The representations are differentiated according to the presence of a LoS propagation path. While both approaches demonstrate satisfactory cluster separation at 10 dB SNR, ContraWiMAE exhibits substantially more distinct class boundaries under degraded signal conditions (0 and 5 dB SNR). Conversely, WiMAE demonstrates considerable misclassification, frequently confounding non-LoS channels with their LoS counterparts under these adverse conditions. This observation underscores the superior noise resilience conferred by the masked contrastive objective.
VI-C Sensitivity to Architectural Design Choices
(a) Top-1 accuracy vs. mask ratio.
(b) Top-1 accuracy vs. model depth.
(c) Top-1 accuracy vs. dimension.
Figure 7: Beam selection performance for WiMAE with linear probing across different model parameters and codebook (CDB) sizes.
To analyze the sensitivity of the model behavior to architecture design choices, we explore different masking ratios $m_{\text{r}}\in\{0.45,0.60,0.75,0.90\}$, network depths $L_{\text{enc}}\in\{6,12,18\}$ and $L_{\text{dec}}\in\{4,8,12\}$, encoding dimensions $d_{\text{e}}\in\{32,64,96,128\}$, and patch configurations $(N_{\text{p,s}},N_{\text{p,f}})\in\{(16,1),(1,16),(4,4)\}$, with fixed attention heads $M_{\text{enc}}=16$ and $M_{\text{dec}}=8$. For example, in Fig. 7 that summarizes the sensitivity results, we use a different patch size $(N_{\text{p,s}},N_{\text{p,f}})$ = (1, 16). We begin by analyzing the impact of the masking ratio, the depth of the model, and the representation dimension on the performance of the WiMAE, providing insights into the sensitivity to design choices for reconstruction-only foundation. To this end, we assess the beam selection task performance along with the reconstruction loss from Eq. (1), which we minimize during pretraining. The architectural choices are critical as they determine the trade-off between computational efficiency and representation quality, which are essential considerations for practical deployment in resource-constrained wireless systems. While our sensitivity analysis focuses on the specific tasks and datasets considered in this work, it provides practical insights into the mechanistic characteristics of our method and guidance for architectural design decisions.
@@ VI-C Sensitivity to Architectural Design Choices
Fig. 7(a) reveals the interplay between mask ratio and feature quality, reflecting the need to balance task difficulty; a small ratio makes reconstruction trivial while a large ratio may render the learning intractable. This suggests an optimal information density for the beam selection task where sufficient masking forces the model to learn meaningful representations without compromising its ability to capture essential discriminative channel characteristics. Moreover, we observe that the same phenomenon applies to reconstruction tasks where the performance is highly dependent on finding a suitable mask ratio. Note that task difficulty is not only determined by the mask ratio but also by patch shape, which plays an important role. We found $m_{\text{r}}=0.6$ to be a fair trade-off point for a patch shape of $(N_{\text{p,s}},N_{\text{p,f}})=(1,16)$.
@@
Fig. 7(b) demonstrates the complex relationship between model depth and performance for the task and dataset considered. As the encoder depth increases from 6 to 12 layers, we observe consistent performance improvements across all codebook sizes. However, the marginal benefit diminishes beyond 12 layers, with 18-layer models showing minimal improvement. This suggests that moderate encoder depth effectively captures channel characteristics without excessive computation or potential overfitting. More interestingly, for a fixed 12-layer encoder, increasing the decoder depth from 4 to 8 and then to 12 layers provides substantial performance improvements. When the decoder is discarded after pretraining, its added complexity does not affect inference latency. This asymmetric architecture—moderate encoder with deep decoder—reveals that a sophisticated decoding process may provide a stronger learning signal that helps the encoder capture more nuanced channel features during training.
@@
Fig. 7(c) reveals a strong positive correlation between embedding dimension and performance, with patterns that vary by task complexity. Top-1 accuracy improves as the dimension increases from 32 to 128, with diminishing returns becoming apparent beyond 96. The most significant improvement occurs between dimensions 32 and 64, suggesting that this range is critical for adequately representing channel characteristics for the data under consideration. More challenging tasks show less saturation at higher dimensions, indicating that complex discrimination tasks may benefit more from increased representational capacity.
VII Conclusion
In this paper, we introduced ContraWiMAE, a transformer-based contrastive masked autoencoder, specifically designed for wireless channel representation learning. By employing an asymmetric encoder-decoder architecture with high masking ratios and moderate encoder depth, ContraWiMAE strikes a balance between representational quality and computational efficiency. The key innovation lies in our novel masked contrastive objective exploiting the inherent characteristics of wireless propagation by applying distinct random masks to noisy observations of identical channel realizations. This approach enables simultaneous extraction of both coherent structural patterns and discriminative characteristics.
Our experiments demonstrated that ContraWiMAE consistently outperforms existing baselines across multiple downstream tasks, achieving particularly strong results under constrained data conditions, a capability that is especially critical for practical wireless scenarios where labeled data are expensive or difficult to obtain. Through extensive evaluation on realistic, unseen wireless scenarios, ContraWiMAE exhibited strong cross-frequency and cross-environment transferability, as well as noise robustness. We showed that, unlike other models that require complex downstream architectures to achieve good performance, ContraWiMAE produces representations that remain effective even with lightweight classifiers. We validated our findings across tasks requiring both discriminative understanding (LoS detection and beam selection) and reconstruction capability (channel estimation). Under challenging low-SNR conditions and distribution shifts, ContraWiMAE matches or surpasses supervised baselines, validating robustness of our wireless-inspired augmentation strategy.
Beyond technical contributions, this work contributes to a paradigm shift toward unified wireless AI systems where a single foundation model serves multiple applications with minimal adaptation overhead. Future work could extend this framework to MIMO configurations, investigate its use for other downstream tasks, and develop efficient compression techniques for edge deployment. Another extension is to include mobility and Doppler effects by analyzing time-domain correlations to capture temporal channel dynamics. While our current evaluation uses ray-tracing data with idealized antenna patterns, future extensions could incorporate hardware impairments and validation on over-the-air measurements to further demonstrate the framework’s robustness in operational deployments. We envision ContraWiMAE serving as a strong baseline, providing a principled approach to self-supervised wireless channel representation learning that respects the unique characteristics of wireless propagation while achieving practical performance across diverse applications. We release the ContraWiMAE model weights and training pipeline to support reproducibility and future advances in self-supervised wireless channel representation learning.